OpenStudy (anonymous):

Solve 2 log2 2 + 2 log2 6 − log2 3x = 3. x = 2 x = 6 x = 16 x = 18

OpenStudy (anonymous):

@IMStuck

OpenStudy (imstuck):

Let me ask you something first...how do you tag someone for a question? Like up above, you have my name with the @ symbol...how do you do that? I have no clue. Tell me and I'll help you! wink wink

OpenStudy (imstuck):

Ok, step by step on this one, ok?

OpenStudy (imstuck):

\[2\log _{2}2+2\log _{2}6-\log _{2}3x=3\]\[\log _{2}(2^{2}*6^{2})-\log _{2}3x=3\]\[\log _{2}(4*36)-\log _{2}3x=3\]\[\log _{2}(144)-\log _{2}3x=3\]

OpenStudy (imstuck):

Look closely step by step. I am doing it one thing at a time, not skipping any steps at all.

OpenStudy (anonymous):

you hold down the shift key then hit the @ symbol and type in whoever you want to tag

OpenStudy (imstuck):

\[\log _{2}\frac{ 144 }{ 3x }=3\]Our rule of logs tells us then that:\[2^{3}=\frac{ 144 }{ 3x }\]\[8=\frac{ 144 }{ 3x }\]Multiply both sides by 3x to get this:\[8(3x)=144\]\[24x=144\]\[x=6\]

OpenStudy (imstuck):

Thank you for that tip! I have been wondering about that forever!

OpenStudy (anonymous):

no, thank you.

OpenStudy (imstuck):

Thanks for the medal!

OpenStudy (imstuck):

ps. logs stink. Just try and memorize the rules for them, ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have one more and then i'll leave you alone

OpenStudy (imstuck):

Let's see it

OpenStudy (anonymous):

Evaluate log6 14. 1.47 0.68 2.33 0.43

OpenStudy (imstuck):

Let me think a sec...rules, rules, rules...

OpenStudy (imstuck):

Ok, here we go...rules rules rules...

OpenStudy (imstuck):

First thing to do is to set the dumb thing = to x. Like this, which will then give us something to solve for.

OpenStudy (imstuck):